# Dimension reduction via score ratio matching

### v1

@@ 7 Conclusions and future work
While the posterior approximation error guarantees in this work are currently limited to reduced subspaces of the parameters and observations, future work will investigate extensions to nonlinear dimension reduction; see Bigoni et al. (2022) for related work in a function approximation setting where gradients are available. In addition, it will be valuable to understand how the intrinsic dimension of the inference problem affects the number of samples needed to learn the score ratio network and to uncover this low-dimensional structure.
@@ A.1 Log-Sobolev inequalities
We start by stating a result achieved in the proof of Corollary 2.11 of Zahm et al. [2022]. Let $U=[U_{r}\ U_{\perp}]\in\mathbb{R}^{n\times n}$ be a unitary matrix, and define $P_{\perp}=U_{\perp}U_{\perp}^{\top}.$ Let $w(x,y)=\nabla_{x}\log\left(\pi_{X|Y}(x|y)/\rho(x)\right)$. Then
@@
$W_{r}W_{r}^{\top}A=\sum_{i=1}^{d}\lambda_{i}(W_{r}W_{r}^{\top})w_{i}w_{i}^{\top}=\sum_{i=1}^{d}\lambda_{i}W_{r}(W_{r}^{\top}w_{i})w_{i}^{\top}=\sum_{i=1}^{r}\lambda_{i}w_{i}w_{i}^{\top}$
@@
$PA=A-\sum_{i=1}^{r}\lambda_{i}w_{i}w_{i}^{\top}=\sum_{i=r+1}^{d}\lambda_{i}w_{i}w_{i}^{\top}$
and $PAP=\sum_{i=r+1}^{d}\lambda_{i}w_{i}w_{i}^{\top}$ follows similarly.
@@
Appendix B Additional details and figures for the numerical results
@@ B.1 Score ratio network hyperparameters
For each problem we use the Adam [Kingma and Ba, 2015] optimizer with a learning rate $5\times 10^{-3}$ and batch size $1000$ to train the network. We take the nuclear norm regularization parameters to be $\lambda_{1}=1/n$ and $\lambda_{2}=1/m$, where $n$ and $m$ are the parameter and observation dimensions. As discussed in Song et al. [2020], directly evaluating the trace operator in the objective function $F$ is prohibitively expensive for even moderate dimensions $n$, and so we also make use of the sliced-score matching method proposed in that work with $100$ projections. The network $\psi_{\theta}$ were taking to be fully connected. Table 1 summarizes all hyperparameter choices for the score ratio network. These parameters include the network depth (i.e., number of hidden layers), width (i.e., layer dimension), the training set size and number of epochs.
@@ B.2 Approximate inference via measure transport
Our underlying motivation for reducing the parameter and observation dimensions is to improve the performance of approximate sampling methods. One sampling strategy that has recently grown in popularity is parametric measure transport (which encompasses normalizing flows), where one attempts to couple the target distribution $\pi$ (e.g., the posterior distribution of interest) with a tractable reference distribution $\rho$ (in our case the standard Gaussian distribution) using a map $S$ so that $S_{\sharp}\rho=\pi$. Having access to this map permits generating i.i.d. samples from the target distribution by generating reference samples $z^{(n)}\sim\rho$ and evaluating the map at these samples $S^{-1}(z^{(n)})\sim\pi$. We direct an interested reader to Marzouk et al. [2016], Papamakarios et al. [2021] and references therein for a comprehensive review of such methods.
In general, the accuracy of the generated posterior samples depends on (1) the expressivity of the transport map class, and (2) the ability to optimize the transport map. In practice, it is often seen that large training sets and computational effort are required to optimize expressive maps for high dimensional problems. Brennan et al. [2020], Baptista et al. [2022] show that dimension reduction can improve the fidelity of approximate posteriors, especially in the data- and compute-limited regimes.
In the numerical experiments of Sections 6.3 and 6.4, we utilize unconstrained monotonic neural networks (UMNNs) [Wehenkel and Louppe, 2019] implemented within the nflows software package [Durkan et al., 2020] as the underlying transport map class. The map $S\colon\mathbb{R}^{n}\times\mathbb{R}^{\operatorname{dim}(y)}\rightarrow\mathbb{R}^{n}$ for sampling conditional distributions depends on the conditioning variable $y$. For an ($n=1$)-dimensional parameter, the map takes the form
@@
where $f:\mathbb{R}^{\operatorname{dim}(y)}\times\mathbb{R}\times\mathbb{R}^{\operatorname{dim}(\phi)}\to\mathbb{R}_{>0}$ is a strictly positive parametric function and $\beta\in\mathbb{R}$ is a scalar bias term. The function $f$ can be made arbitrarily complex using an unconstrained neural network whose output is forced to be strictly positive through a shifted ELU activation unit. Here $\phi$ denotes the parameters of this neural network, meaning the map parameters we train are $(\phi,\beta)$. For both example, the underlying networks have $5$ hidden layers. We let the number of hidden features scale with the input-dimension, taking it to be equal to $\lceil 1.5\operatorname{dim}(y)\rceil$. We train the map using Adam [Kingma and Ba, 2015] using a step-size of $5\times 10^{-4}$. $10$% of the training samples were held out to be a validation set and training was stopped when the objective evaluated on this validation set plateaus in value.

### v2

@@ 7 Conclusions and future work
We outline several promising directions for future work. While the posterior approximation error guarantees in this work are currently limited to reduced subspaces of the parameters and observations, future work will investigate extensions to nonlinear dimension reduction; see Bigoni et al. (2022) for related work in a function approximation setting where gradients are available. In addition, it will be valuable to understand how the intrinsic dimension of the inference problem, along with the smoothness of the underlying densities, affect the number of samples $N$ needed to estimate the score ratio and to uncover this low-dimensional structure. For example, if one can characterize how the error in an estimated score ratio $\hat{w}_{\theta}^{N}$
$\mathbb{E}_{\pi_{X,Y}}\left\|\hat{w}_{\theta}^{N}(x,y)-\nabla_{x}\log\left(\frac{\pi_{X|Y}(x|y)}{\rho(x)}\right)\right\|^{2},$
in expectation over the sample, scales with an appropriate intrinsic dimension and sample size, then Theorem 3.3 could be reformulated using user-specified parameters. We also note that Theorem 3.3 currently addresses parameter dimension reduction but not observation dimension reduction. Future work will aim to establish an analogous result for observation reduction, thereby mirroring the duality between Proposition 2.2 and Proposition 2.3.
@@ A.1 Log-Sobolev inequalities
We start by stating a result achieved in the proof of Corollary 2.11 of Zahm et al. (2022). Let $U=[U_{r}\ U_{\perp}]\in\mathbb{R}^{n\times n}$ be a unitary matrix, and define $P_{\perp}=U_{\perp}U_{\perp}^{\top}.$ Let $w(x,y)=\nabla_{x}\log\left(\pi_{X|Y}(x|y)/\rho(x)\right)$. Then
@@
$\Phi_{r}\Phi_{r}^{\top}A=\sum_{i=1}^{d}\lambda_{i}(\Phi_{r}\Phi_{r}^{\top})\phi_{r}\phi_{r}^{\top}=\sum_{i=1}^{d}\lambda_{i}\Phi_{r}(\Phi_{r}^{\top}\phi_{r})\phi_{r}^{\top}=\sum_{i=1}^{r}\lambda_{i}\phi_{r}\phi_{r}^{\top}$
@@
$PA=A-\sum_{i=1}^{r}\lambda_{i}\phi_{r}\phi_{r}^{\top}=\sum_{i=r+1}^{d}\lambda_{i}\phi_{r}\phi_{r}^{\top}$
and $PAP=\sum_{i=r+1}^{d}\lambda_{i}\phi_{r}\phi_{r}^{\top}$ follows similarly.
@@
Appendix B Additional details for the numerical results
@@ B.1 Score ratio network hyperparameters
For each problem we use the Adam (Kingma & Ba, 2015) optimizer with a learning rate $5\times 10^{-3}$ and batch size $1000$ to train the network. We take the nuclear norm regularization parameters to be $\lambda_{1}=1/n$ and $\lambda_{2}=1/m$, where $n$ and $m$ are the parameter and observation dimensions. As discussed in Song et al. (2020), directly evaluating the trace operator in the objective function $F$ is prohibitively expensive for even moderate dimensions $n$, and so we also make use of the sliced-score matching method proposed in that work with $100$ projections. The network $\psi_{\theta}$ were taking to be fully connected. Table 1 summarizes all hyperparameter choices for the score ratio network. These parameters include the network depth (i.e., number of hidden layers), width (i.e., layer dimension), the training set size and number of epochs.
@@ B.2 Approximate inference via measure transport
Our underlying motivation for reducing the parameter and observation dimensions is to improve the performance of approximate sampling methods. One sampling strategy that has recently grown in popularity is parametric measure transport (which encompasses normalizing flows), where one attempts to couple the target distribution $\pi$ (e.g., the posterior distribution of interest) with a tractable reference distribution $\rho$ (in our case the standard Gaussian distribution) using a map $S$ so that $S_{\sharp}\rho=\pi$. Having access to this map permits generating i.i.d. samples from the target distribution by generating reference samples $z^{(n)}\sim\rho$ and evaluating the map at these samples $S^{-1}(z^{(n)})\sim\pi$. We direct an interested reader to Marzouk et al. (2016); Papamakarios et al. (2021) and references therein for a comprehensive review of such methods.
In general, the accuracy of the generated posterior samples depends on (1) the expressivity of the transport map class, and (2) the ability to optimize the transport map. In practice, it is often seen that large training sets and computational effort are required to optimize expressive maps for high dimensional problems. Brennan et al. (2020); Baptista et al. (2022) show that dimension reduction can improve the fidelity of approximate posteriors, especially in the data- and compute-limited regimes.
In the numerical experiments of Sections 6.3 and 6.4, we utilize unconstrained monotonic neural networks (UMNNs) (Wehenkel & Louppe, 2019) implemented within the nflows software package (Durkan et al., 2020) as the underlying transport map class. The map $S\colon\mathbb{R}^{n}\times\mathbb{R}^{\operatorname{dim}(y)}\rightarrow\mathbb{R}^{n}$ for sampling conditional distributions depends on the conditioning variable $y$. For an ($n=1$)-dimensional parameter, the map takes the form
@@
where $f:\mathbb{R}^{\operatorname{dim}(y)}\times\mathbb{R}\times\mathbb{R}^{\operatorname{dim}(\phi)}\to\mathbb{R}_{>0}$ is a strictly positive parametric function and $\beta\in\mathbb{R}$ is a scalar bias term. The function $f$ can be made arbitrarily complex using an unconstrained neural network whose output is forced to be strictly positive through a shifted ELU activation unit. Here $\phi$ denotes the parameters of this neural network, meaning the map parameters we train are $(\phi,\beta)$. For both example, the underlying networks have $5$ hidden layers. We let the number of hidden features scale with the input-dimension, taking it to be equal to $\lceil 1.5\operatorname{dim}(y)\rceil$. We train the map using Adam (Kingma & Ba, 2015) using a step-size of $5\times 10^{-4}$. $10$% of the training samples were held out to be a validation set and training was stopped when the objective evaluated on this validation set plateaus in value.
B.3 Derivation of the posterior and score function for the Flux problem
The flux is defined as
where $\Gamma_{\text{B}}=[0,1]\times\{0\}$ denotes the bottom boundary of the domain. As done section 6.2, we express the log-diffusivity field $\varrho$ in the Karhunen-Loève expansion of the Gaussian prior with coefficients $x$, defining the mapping $\mathcal{Q}:\mathbb{R}^{100}\to\mathbb{R}$
$q=\mathcal{Q}(x).$
The target posterior for this problem, $\pi_{Q|Y}$, can then be written as the following integral via change of measure formula
$\pi_{Q|Y}(q|y)=\int_{{\mathcal{Q}^{-1}(q)}}\frac{\pi_{X|Y}(x|y)}{\left\|\nabla_{x}\mathcal{Q}(x)\right||}{\rm d}x.$
We note here that evaluating the unnormalized posterior and thus the posterior score function is intractable, as it requires the computing the integral over the pre-image of the mapping $\mathcal{Q}$.